OpenStudy (anonymous):

solve the following systems of equations and state the coorindates of all points in the solution set. y=x^2+4x-5 y=x-1 ( ) , ( ) and ( ) , ( )

OpenStudy (anonymous):

@brinethery

OpenStudy (anonymous):

It's asking for the places where y=0, correct? B/c that's how I'll solve it.

OpenStudy (anonymous):

x-1 = x^2 +4x -5 x^2 +3x -4 = 0 (x+4)(x-1) = 0 When y = 0, then x=-4 and x=1 Your points: (-4,0) and (1,0)

OpenStudy (anonymous):

@Helllo;3 are you there???